OpenStudy (anonymous):

P(8, 5) = 40,320 1,680 6,720

OpenStudy (anonymous):

More information?

OpenStudy (anonymous):

That is it. That is the question! :/

OpenStudy (anonymous):

What were you learning?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8p5 = \Large {8! \over (8 - 5)!} = 8 \times 7\times 6\times 5 \times 4 }\)

OpenStudy (anonymous):

Ohhh. nPr.

Parth (parthkohli):

Haha you thought of probability :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow nPr = \Large{n! \over (n - r)!} }\)

OpenStudy (anonymous):

lol Yeah. You beat me agin. lol I was going to put the formula :)

Parth (parthkohli):

\(\color{Red}{\Rightarrow 8P5 = \Large {8! \over (8 - 5)!} = {8! \over 3!} = {8 \times 7 \times 6 \times 5\times 4 \times \cancel{3\times 2 \times 1} \over \cancel{3 \times 2 \times 1}} = 8 \times 7 \times 6 \times 5\times 4 }\)